OpenStudy (anonymous):

Simlify 27 - 3x / 5x- 45. - Can someone please get me through this step by step ??

OpenStudy (anonymous):

\[\frac{ 3(9-x) }{ 5(x - 9) }\] \[\frac{ 3 }{ 5 } \times \frac{ -1(x - 9) }{( x-9) }\] Can you solve it now?

OpenStudy (anonymous):

Muchas Gracias ( : !

OpenStudy (anonymous):

Haha, sure! :D